OpenStudy (anonymous):

What is the largest possible remainder when a two-digit number is divided by the sum of its digits? a) 13 b) 14 c) 15 d) 16 e) 17

OpenStudy (anonymous):

interesting problem. do you know the answer/solution?

OpenStudy (anonymous):

No I do not know the solution :(

sam (.sam.):

The largest possible digit sum = 9 + 9 = 18, so no remainder can be > 17. Working down from the top: 99/18 → remainder 9 98/17 → remainder 13 89/17 → remainder 4 97/16 → remainder 1 88/16 → remainder 8 79/16 → remainder 15 After that the divisor is 15 or less, so the remainder is 14 or less, and so the largest remainder is 15, from 79 = 4 * 16 + 15.

OpenStudy (anonymous):

wow.

OpenStudy (anonymous):

that was quick...

OpenStudy (anonymous):

Thank you! :D